# Submitted to Phys. Rev. B. Excitonic Insulator and the Extended Falicov–Kimball Model Away from Half-Filling

### v1

@@ VI DISCUSSION
The possibility of excitonic condensation, resulting in a formation of EI state in a metallic or semiconducting compound at a low temperature, attracts broad experimental and theoretical effort (see Ref. 26 for a contemporary review). This includes theoretical studies of doped systems, with applications to hexaborides[27, 16, 28, 29], twisted bilayer graphene[17], and other compounds[18]. These broader-band systems are generally treated within the low-energy, long-wavelength approach, which in its original form involves re-structuring of the spectrum in the immediate vicinity of the (nested) Fermi surfaces, and lowering the net energy slightly by opening a narrow gap at the Fermi level[30]. When doped, the proximity of the EI state affects the properties of the system only as long as the chemical potential lies close to the (possibly smeared) excitonic gap, corresponding to a narrow range of doping values. Within this range, a rich physical picture emerges once additional degrees of freedom (most notably spin) and features like imperfect nesting[27, 16, 17] or presence of impurities[28] are taken into account. Findings include, inter alia, ferromagnetism[31, 32, 27, 16, 28, 29], excitonic metal behaviour[16, 18, 17, 32], and phase separation[27, 16, 29].
@@
[…]

### v2

@@ VI DISCUSSION
The possibility of excitonic condensation, resulting in a formation of EI state in a metallic or semiconducting compound at a low temperature, attracts broad experimental and theoretical effort (see Ref. 26 for a contemporary review). This includes theoretical studies of doped systems, with applications to hexaborides[27, 18, 28, 29], twisted bilayer graphene[19], and other compounds[20]. These broader-band systems are generally treated within the low-energy, long-wavelength approach, which in its original form involves re-structuring of the spectrum in the immediate vicinity of the (nested) Fermi surfaces, and lowering the net energy slightly by opening a narrow gap at the Fermi level[30]. When doped, the proximity of the EI state affects the properties of the system only as long as the chemical potential lies close to the (possibly smeared) excitonic gap, corresponding to a narrow range of doping values (cf. Appendix). Within this range, a rich physical picture emerges once additional degrees of freedom (most notably spin) and features like imperfect nesting[27, 18, 19] or presence of impurities[28] are taken into account. Findings include, inter alia, ferromagnetism[31, 32, 27, 18, 28, 29], excitonic metal behaviour[18, 20, 19, 32], and phase separation[27, 18, 29].
@@
Either way, we expect that EI behaviour, which is conventionally associated with EFKM at half-filling, may affect the properties of a doped system over a broad range of carrier densities. Therefore, whenever experimental results suggest that the EI or EM behaviour persists beyond one or two percent doping, this may imply that the EFKM-like picture of strong short-range correlations is relevant. In particular, this might be the case for 1TT-TiSe2 (where the charge density wave, broadly attributed to excitonic pairing, is cut off by a superconducting state at 4%-6% copper intercalation[37]), and likely also for Ta2NiSe5 (see Refs. 38, 39).
@@ Acknowledgements.
Appendix A Compressibility of the excitonic metal phase
At the phenomenological level, the origin of the negative compressibility in the EM phase away from half-filling is clear from Fig. 1. Both above and below the half-filling the EM eventually merges with another phase, and the difference between the values of $\mu$ at these two points, whether negative (as in Fig. 1 b) or positive (Fig. 1 a), is relatively small. However, the dependence of $\mu(n)$ in the intervening EM phase must accommodate a large positive jump at $n=1$, equal to the width of the EI energy gap. It is therefore quite natural that away from this jump the value of $\mu$ decreases with $n$.
In order to analyse the underlying mechanism in more detail, we will turn to the $t^{\prime}\rightarrow 0$ case, ignoring the other (collective-excitation) instability[7] which arises there. This is expedient because the negative compressibility of EM persists for $t^{\prime}\rightarrow 0$ as well and clearly has the same nature as for $t^{\prime}\neq 0$, while on the other hand the $t^{\prime}=0$ case is much more transparent analytically.
At $t^{\prime}=0$, the quasiparticle energies $\epsilon^{1,2}_{\vec{k}}=\epsilon_{1,2}(\epsilon_{\vec{k}})$, Eqs. (67), are monotonously increasing functions of the tight-binding energy $\epsilon_{\vec{k}}$. Thus for $n<1$ ($n>1$) the quasiparticle states in the lower (upper) band are occupied for $\epsilon_{\vec{k}}<\epsilon_{0}$, and the chemical potential $\mu$ is given by $\epsilon_{1}(\epsilon_{0})$ $\epsilon_{2}(\epsilon_{0})$ for $n>1$]. When the carrier density is increased by a small $\delta n$, the chemical potential change $\delta\mu=\delta_{\rm r}\mu+\delta_{\rm c}\mu$ is obtained from Eq.(6) as a sum of the “rigid” bandstructure contribution,
$\delta_{\rm r}\mu=\frac{1}{2}\left(1\pm\frac{\xi_{0}}{\sqrt{\xi_{0}^{2}+4U^{2}\Delta^{2}}}\right)\delta\epsilon_{0}$ (21)
[with $\delta\epsilon_{0}=\delta n/\nu(\epsilon_{0})$ and $\xi_{0}=\xi(\epsilon_{0})$, see Eq. (7)], which is always positive, and the interaction-induced “correlation” term,
$\delta_{\rm c}\mu=\frac{U\delta n}{2}\mp\frac{\xi_{0}U(\delta n_{c}-\delta n_{d})+4U^{2}\Delta\delta\Delta}{2\sqrt{\xi_{0}^{2}+4U^{2}\Delta^{2}}}\,$, (22)
which includes the effects of a self-consistent change of $n_{c,d}$ and $\Delta$. The upper and lower signs in Eqs. (2122) correspond to $n<1$ and $n>1$ respectively.
By varying $n$ in the mean field equations (34) and solving for $\delta n_{c}-\delta n_{d}$ and $\delta\Delta$, we find that $\delta_{\rm c}\mu$ is proportional to $\delta n$,
$\delta_{\rm c}\mu$ $=$ $\frac{(2\xi_{0}Y_{0}-2U\Delta Y_{1}-U\Delta)^{2}+(4U\Delta Y_{0}+\xi_{0}Y_{1})^{2}}{16Y_{0}^{2}+4Y_{1}^{2}+2Y_{1}}\times$ (23)
$\times\frac{2U\delta n}{\xi_{0}^{2}+4U^{2}\Delta^{2}}\,$,
where [cf. Ref. 7, Eqs.(B3-B4)]
$Y_{0}=\frac{1}{2}\int\frac{U^{2}\xi\Delta\nu(\epsilon)d\epsilon}{(\xi^{2}+4U^{2}\Delta^{2})^{3/2}}\,,\,\,\,Y_{1}=-\int\frac{2U^{3}\Delta^{2}\nu(\epsilon)d\epsilon}{(\xi^{2}+4U^{2}\Delta^{2})^{3/2}}\,$, (24)
$\xi=\xi(\epsilon)$, the integration limits are from d-d to $\epsilon_{0}$ for $n<1$ or from $\epsilon_{0}$ to $d$ for $n>1$, and $d=2$ or $3$ is the dimensionality.
The denominator in Eq. (23) is negative, as proved in Ref. 7 (Appendix B). The proof, based on the required stability of the EI state at a fixed $n=1$ with respect to an “external field” $V_{0}$ [see Eq. (13)], can be directly generalised for the case of an EM state away from half-filling.
The quantity $\delta_{\rm c}\mu/\delta n$ is therefore negative, and gives a negative contribution to $\partial\mu/\partial n$. This has a clear physical meaning: while in the absence of correlations the compressibility would have been given by $\delta_{\rm r}\mu/\delta n>0$, in reality the self-consistent mean field band structure changes with $n$ in such a way that the net energy is minimised. The associated energy reduction is of second order in $\delta n$, which yields a negative term in the compressibility.
The EM phase is found in the region where the quasiparticle density of states at the Fermi level is strongly increased due to hybridisation, meaning $\xi_{0}<0$ for $n<1$, $\xi_{0}>0$ for $n>1$ (i.e., the Fermi level lies between the crossing, $\xi(\epsilon)=0$, of the Hartree $\Delta\rightarrow 0$ bands and the EM energy gap). In the $U\Delta\ll|\xi_{0}|$ limit, the term $\delta_{\rm r}\mu/\delta n$ then vanishes, whereas $Y_{1}$ and hence $\delta_{\rm c}\mu/\delta n$ remain finite. Similarly in the $U\Delta\gg 1$ limit, $\delta_{\rm r}\mu/\delta n$ remains finite whereas $\delta_{\rm c}\mu/\delta n\propto-U\Delta$. Both limits are formal since they can not arise as mean-field solutions; the real $\Delta\rightarrow 0$ regime at $t^{\prime}=0$ corresponds to either $n_{d}\rightarrow 0$ or $n_{d}\rightarrow 1$. Yet the fact that the negative $\delta_{\rm c}\mu$ dominates in both cases, along with the results of our numerical investigation of actual mean field solutions for the 2D case, suggests that this is indeed the generic behaviour, both in $t^{\prime}=0$ and $t^{\prime}\neq 0$ cases.
There is however one important exception, corresponding to the 3D case just above or just below half-filling. In the limit $n\rightarrow 1$, owing to vanishing $\nu(\epsilon)$ at the band edge, the dominant term is $\delta_{\rm r}\mu\propto(\delta n)^{2/3}$. This results in a positive compressibility, presumably within an immediate vicinity of half-filling[40].
Based on the above discussion of EM compressibility in an oversimplified model, we can formulate tentative conclusions which perhaps are not restricted to the EFKM as described by Eq.(1): (i) $\delta\mu$ contains two terms, corresponding to filling a rigid hybridised bandstructure and to correlation effects. The correlations term is negative, always linear in $\delta n$ (because it comes from a variation of Fermi-sea integrals) and typically dominant, resulting in a negative $\partial\mu/\partial n$ (ii) In three dimensions, whenever the Fermi level approaches a band edge or a Van Hove anomaly, the compressibility turns positive, presumably within a narrow range of doping values. Thus the EM may be thermodynamically stable in this region, provided that there are no other instabilities. In this regard, we note that the purported excitonic ferromagnetism in hexaborides is limited to small values of doping, although the difference in both physical systems and theoretical approaches does not allow for a direct analogy (see discussion in Sec. VI).